Jim and Amy sold many pieces of furniture and made $1380 during their garage sale. They had 16 more $10 bills then $50 dollar bills. They had one more then 3 times as many $20 bills as $50 bills. How many of each denomination did they have?
Answer by lwsshak3(11628)

Jim and Amy sold many pieces of furniture and made $1380 during their garage sale. They had 16 more $10 bills than $50 dollar bills. They had one more then 3 times as many $20 bills as $50 bills. How many of each denomination did they have?
***
let x=number of $50 bills
x+16=number of $10 bills
3x+1=number of $20 bills
..
50x+10(x+16)+20(3x+1)=1380
50x+10x+160+60x+20=1380
120x=1380-180=1200
x=10
x+16=26
3x+1=31
number of $50 bills=10
number of $10 bills=26
number of $20 bills=31